# Revolutionizing Large Language Model Training through Dynamic Parameter Adjustment

### v1

@@
Abstract:   In the era of large language models, the demand for efficient use of computational resources has become critically important. Although parameter-efficient fine-tuning techniques have achieved results comparable to full fine-tuning, their application during the pre-training phase poses significant challenges. Specifically, employing parameter-efficient strategies at the onset of pre-training can severely compromise efficiency, especially in larger models. In this paper, building upon the fine-tuning method LoRA, we introduce a novel parameter-efficient training technique that frequently alters trainable part of parameters, facilitating effective pre-training. Our method not only achieves memory reductions and computational overhead comparable to current state-of-the-art parameter-efficient algorithms during the pre-training phase but also maintains accuracy levels comparable to those of full pre-training. We provide both theoretical analyses and empirical evidence to demonstrate the effectiveness of our approach.
@@ 1 Introduction
In recent years, the size of large language models (LLMs) has increased rapidly, a trend ignited by the advent of the transformer architecture [33]. To support the training of increasingly large models, various distributed training techniques have been employed. These include data parallelism [4, 19], the Zero Redundancy Optimizer [27] to optimize data splitting, tensor parallelism [28], and pipeline parallelism [16, 26] to partition the model itself. Despite these advancements, hardware resources are increasingly becoming a bottleneck as model sizes continue to expand [26]. For models exceeding one trillion parameters, the predominant approach is 3D parallelism, which integrates data, tensor, and pipeline parallelism. Presently, a significant portion of the inter-node communication overhead is attributed to data parallelism, required by the frequent communication of large quantities of parameter gradients.
To address these challenges, various parameter-efficient strategies have been proposed. Techniques such as model sparsification [1, 29] and progressive model pruning during training [7] have shown promise. Additionally, methods leveraging Singular Value Decomposition (SVD) to approximate full-rank matrices in low-rank spaces have been explored [30, 34, 38, 39].
Beyond the entire training process, several techniques focus on the fine-tuning phase to enhance adaptability and efficiency. For instance, the Adapter method [14, 13] introduces additional trainable layers while freezing the remaining parameters. Similarly, prefix-tuning [20] employs a small set of trainable vectors (prefixes) at the beginning of each transformer layer to enable adaptation to new tasks without altering the core pre-trained model parameters.
Another noteworthy fine-tuning strategy is Low-Rank Adaptation (LoRA) [15], which has attracted attention for its effectiveness and the fact that it introduces no additional computational overhead during inference. However, it is important to note that while these parameter-efficient methods like LoRA are highly effective for fine-tuning, directly applying them during the pre-training phase can lead to a considerable loss in model accuracy as observed in many prior works [34, 35, 24]. We hypothesize that this loss of accuracy is due to the premature use of low-rank training, which may cause the model to become trapped in a local minimum. In response, we aim to develop a strategy that utilizes parameter-efficient training while frequently changing the trainable part of parameters. This approach is designed to mimic the behavior of full-rank training, potentially overcoming the limitations observed with early implementations of low-rank adaptations.
@@
  • We have developed a novel strategy utilizing Low-Rank Adaptation (LoRA), termed SwiLoRA, which enables smooth adjustments to the trainable parameters of the LoRA matrices, as illustrated in Figure 1. This approach extends the application of LoRA to the pre-training stage with minimal loss of accuracy. Additionally, SwiLoRA is scalable to any model size, as it introduces negligible additional computational overhead.
  • We propose a method for calculating optimal initial parameters for LoRA matrices and their associated candidate vectors introduced in our algorithm. This initiative is designed to accelerate the warm-up phase of training and ensure balanced updates of the LoRA matrices, thereby improving the overall efficiency of the training process.
  • We conduct a theoretical analysis of the properties of the local linear system incorporating SwiLoRA, providing insights into how SwiLoRA affect the system’s dynamics and performance.
@@ 2 Methodology
A substantial body of research, such as various pruning methods [10, 3], has demonstrated that neural networks tend to exhibit low-rank characteristics after certain stages of training. Techniques for parameter-efficient fine-tuning, such as LoRA, capitalize on this observation. Concurrently, studies like [23, 9] have revealed that overparameterization in neural networks can lead to implicit regularization, thereby enhancing generalization. These findings underscore the importance of training with full parameters during the initial phase. Further empirical evidence supporting this phenomenon is provided in works like [34, 35, 24, 39]. Based on these insights, this section proposes a method designed to train a substantial number of parameters, while selectively updating only a portion of the parameters at any one time to conserve computational resources.
Low-Rank Adaptation (LoRA)
@@
Consider a pre-trained model with a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ from a specific linear layer. LoRA proposes an innovative modification: transforming $\mathbf{W}$ into $\mathbf{W}+\mathbf{B}\mathbf{A}$. Here, $\mathbf{B}\in\mathbb{R}^{m\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times n}$ are newly introduced matrices, where $r$ is significantly smaller than both $m$ and $n$. Then during fine-tuning, $\mathbf{W}$ is kept frozen while matrices $\mathbf{B}$ and $\mathbf{A}$ are trained. At the inference stage, $\mathbf{B}\mathbf{A}$ is added to $\mathbf{W}$ which preserves the model’s original structure. The matrix $\mathbf{A}$ is initialized using Kaiming initialization [12], while $\mathbf{B}$ is initially set to a zero matrix to ensure consistency.
SwiLoRA training process
Now, let us delve deeper into the linear system $(\mathbf{W}+\mathbf{B}\mathbf{A})\mathbf{x}=\mathbf{y}$. As illustrated in Figure 1, we decompose the matrix $\mathbf{B}$ into its column vectors $\mathbf{b}_{k}\in\mathbb{R}^{m\times 1}$ for $k=1,\ldots,r$, represented as $\mathbf{B}=[\mathbf{b}_{1},\ldots,\mathbf{b}_{r}]$. Similarly, we decompose matrix $\mathbf{A}$ into its row vectors $\mathbf{a}_{k}^{T}\in\mathbb{R}^{n\times 1}$ for $k=1,\ldots,r$, leading to $\mathbf{A}^{T}=[\mathbf{a}_{1}^{T},\ldots,\mathbf{a}_{r}^{T}]$. Hereafter, we call these vectors $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ as LoRA vectors.
@@
It is known for $k$ matrices $\mathbf{W}_{1},\mathbf{W}_{2},\ldots,\mathbf{W}_{k}$, there holds
@@
If we adopt the strategy in LoRA to add $\mathbf{B}\mathbf{A}$ to $\mathbf{W}$ and only update $\mathbf{B}$ and $\mathbf{A}$ from the pre-training stage, according to (2), the rank of updated parameters of the local linear system through the entire training process will be limited to $2r$. This limitation can potentially impede the training efficacy. To mitigate this issue, we alter the values of $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ to $\mathbf{b}^{\prime}_{k}\in\mathcal{C}(\mathbf{B})$ and $\mathbf{a}^{\prime}_{k}\in\mathcal{C}(\mathbf{A}^{T})$ at appropriate frequencies, respectively, with these new values randomly selected from predefined candidate vectors list $\mathcal{C}(\mathbf{B})$ and $\mathcal{C}(\mathbf{A}^{T})$(one of $\mathbf{b}_{k}^{\prime}$ or $\mathbf{a}_{k}^{\prime}$ can be the same as $\mathbf{b}_{k}$ or $\mathbf{a}_{k}$). To maintain the consistency of the model’s output, compensatory adjustments are made to $\mathbf{W}$. To be more precise, when $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ are updated to $\mathbf{b}^{\prime}_{k}$ and $\mathbf{a}^{\prime}_{k}$, we accordingly adjust $\mathbf{W}$ with the equation $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime T}$.
@@
When implementing these updates, the updated parameters of both $\mathbf{B}$ and $\mathbf{A}$ are derived from $\min(m,n)$ distinct candidate vectors, which ensures updated parameters are full-rank. Readers can refer to [24, 41] for more details.
@@
As previously mentioned, the model initially possesses full rank, and the rank of each layer decreases progressively over time. Consequently, we have adopted an exponential decay function for the switching frequency, where the coefficients are determined empirically. The selection of LoRA rank $r$ for $\mathbf{B}\mathbf{A}$ is influenced by the final rank of the layers post-training, which has been extensively explored in LoRA and its variants [32, 37].
@@
Currently Large Language Models (LLMs) predominantly utilize Adam [18] and AdamW [25] optimizers over SGD, which rely on optimizer states. It is crucial to note that after switching LoRA vectors, the gradients associated with these parameters are also changed, which prevents the reuse of optimizer states. To address this issue, when $\mathbf{a}_{k}$ is switched, we reset the optimizer states of $\mathbf{b}_{k}$. And conversely, when $\mathbf{b}_{k}$ is switched, we reset optimizer states of $\mathbf{a}_{k}$. Note that we reset optimizer states of counterpart pair rather than optimizer states of the switched parameters itself. This approach will be further explained in Section 3. Additionally, when the optimizer states are reset to zero, we freeze corresponding parameters for $N$ steps to maintain the robustness of the training. In this study, $N$ is set to $5$.
The above process is described in Algorithm 1 and Algorithm 2.
@@
Algorithm 2 SwiLoRA training process. switch_num$(step,r,interval_{0},\theta)$ is an integer generator function which yields $\lfloor s\rfloor+X$ numbers sampled from $1$ to $r$ where $s=r/(interval_{0}e^{\theta step})$ and random variable $X\sim\text{Bernoulli}(s-\lfloor s\rfloor)$, i.e. $P(X=1)=1-P(X=0)=s-\lfloor s\rfloor$.
@@
4:        for $i$ in switch_num$(step,r,interval_{0},\theta)$ do
5:           Sample $j\sim\{k\}_{k=1}^{\min(m,n)}$
6:           $\mathbf{W},\mathbf{B},\mathbf{A}\leftarrow\text{switch}(\mathbf{W},\mathbf{B},\mathbf{A},i,j)$
7:           Freeze $\mathbf{A}_{i,:}$ for $N$ steps
8:        end for
9:        for $i$ in switch_num$(step,r,interval_{0},\theta)$ do
10:           Sample $j\sim\{k\}_{k=1}^{\min(m,n)}$
11:           $\mathbf{W}^{T},\mathbf{A}^{T},\mathbf{B}^{T}\leftarrow\text{switch}(\mathbf{W}^{T},\mathbf{A}^{T},\mathbf{B}^{T},i,j)$
12:           Freeze $\mathbf{B}_{:,i}$ for $N$ steps
13:        end for
@@
Initialization of SwiLoRA
Results in [11, 36] have demonstrated the importance of initialization of LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ to the training effects. Unlike these works, which are applied only during the fine-tuning stage, our method is utilized throughout the entire training process. Moreover, zero initialization for $\mathcal{C}(\mathbf{B})$ encounters similar issues when switching candidates. Hence, we cannot simply use the same initialization strategy for $\mathbf{B}$ and $\mathbf{A}$ as employed in LoRA. To achieve appropriate initialization for matrices $\mathbf{B}$ and $\mathbf{A}$ along with their candidate vectors, we follow the idea of Xavier initialization [8] and Kaiming initialization [12].
The main idea of [8] and [12] is to is to maintain a balance in the variance of the activations and gradients across layers during forward and backward propagation. In this study, we focus on balancing the variance of activations. Furthermore, we aim to ensure the updated parameters derived from $\mathbf{B}$ are of the same amount as those derived from $\mathbf{A}$:
$\Delta\mathbf{B}\mathbf{A}=\mathbf{B}\Delta\mathbf{A}.$ (3)
In this paper, we fix the scaling factor of $\mathbf{B}\mathbf{A}$ to $\frac{1}{r}$. Then the weight matrix of the linear layer is modified to $\mathbf{W}+\frac{1}{r}\mathbf{B}\mathbf{A}$. To meet the above conditions, the standard deviation of initialization for $\mathbf{B}$ and $\mathbf{A}$ and their LoRA vectors are

### v2

# SwitchLoRA: Switched Low-Rank Adaptation Can Learn Full-Rank Information

@@
Abstract
In the training of large language models, parameter-efficient techniques such as LoRA optimize memory usage and reduce communication overhead during the fine-tuning phase. However, applying such techniques directly during the pre-training phase results in poor performance, primarily because the premature implementation of low-rank training significantly reduces model accuracy. Existing methods like ReLoRA and GaLore have attempted to address this challenge by updating the low-rank subspace. However, they still fall short of achieving the accuracy of full-rank training because they must limit the update frequency to maintain optimizer state consistency, hindering their ability to closely approximate full-rank training behavior. In this paper, we introduce SwitchLoRA, a parameter-efficient training technique that frequently and smoothly replaces the trainable parameters of LoRA adapters with alternative parameters. SwitchLoRA updates the low-rank subspace incrementally, targeting only a few dimensions at a time to minimize the impact on optimizer states. This allows a higher update frequency, thereby enhancing accuracy by enabling the updated parameters to more closely mimic full-rank behavior during the pre-training phase. Our results demonstrate that SwitchLoRA actually surpasses full-rank training, reducing perplexity from 15.23 to 15.01 on the LLaMA 1.3B model while reducing communication overhead by 54% on the LLaMA 1.3B model. Furthermore, after full fine-tuning the SwitchLoRA pre-trained model and the full-rank pre-trained model on the GLUE benchmark, the SwitchLoRA pre-trained model showed an average accuracy gain of about 1% over the full-rank pre-trained model. This demonstrates enhanced generalization and reasoning capabilities of SwitchLoRA.
@@ 1 Introduction
The size of large language models (LLMs) has increased rapidly due to the advent of the transformer architecture [47]. To support the training of large models, distributed training techniques such as data parallelism [5, 29], tensor parallelism [41], pipeline parallelism [22, 39] and the Zero Redundancy Optimizer [40] have been employed. However, distributed training of trillion-scale models incurs significant inter-node communication overhead from synchronizing extensive parameter gradients across multiple nodes.
To address these challenges, various parameter-efficient strategies have been proposed. Techniques such as model sparsification [1, 42] and progressive model pruning during training [11] have shown promise. Additionally, methods leveraging Singular Value Decomposition (SVD) to approximate full-rank matrices in low-rank spaces have been explored [43, 49, 59]. Beyond the entire training process, several techniques improve adaptability and efficiency during the fine-tuning phase. For example, methods such as the Adapter [20, 18] and Prefix-tuning [30] introduce additional trainable layers while freezing the remaining parameters.
Another noteworthy fine-tuning strategy is Low-Rank Adaptation (LoRA) [21], which introduces no computational overhead during inference while maintaining training accuracy. However, previous studies [49, 50, 34] have observed that parameter-efficient methods such as LoRA perform less efficiently during the pre-training phase because the premature use of low-rank training leads to a considerable loss in model accuracy. To increase the rank of updated parameters and benefit from low-rank training, ReLoRA [34] applies the structure of LoRA and periodically resets LoRA adapters. Similarly, GaLore [60] projects gradients onto a subspace, updating this subspace periodically. These approaches update the descent direction of trainable parameters to mimic the behavior of full-rank training, thereby overcoming the limitations observed in existing implementations of low-rank adaptation. However, we find that the intervals between resetting/updating steps in ReLoRA and GaLore are set to relatively large values because too frequent changes in the updating direction can cause inconsistency in optimizer states, which may not sufficiently approximate the behavior of full-rank training, resulting in a loss of accuracy.
To address this challenge, as illustrated in Figure 1, we introduce SwitchLoRA, which enables smooth and frequent adjustments to the trainable parameters of the LoRA matrices while introducing negligible additional computational overhead. SwitchLoRA maintains a set of candidate vectors for each matrix within the LoRA adapters. At each training step, it replaces portions of the column or row vectors with these candidate vectors, subsequently training the LoRA adapters. This process minimizes the impact on optimizer states, thus allowing for a higher update frequency compared to ReLoRA and GaLore. By more closely approximating full-rank parameter updating behaviors during the pre-training phase, this approach enhances overall accuracy.
@@
  • We propose SwitchLoRA to facilitate smooth and frequent adjustments to the trainable parameters of the LoRA matrices through low-rank adaptation, maintaining the accuracy of full-rank training while reducing communication overhead.
  • To mitigate inconsistencies in optimizer states when parameters are switched, SwitchLoRA resets the corresponding optimizer states and temporarily freezes the affected parameters. Additionally, SwitchLoRA employs a different initialization rule for LoRA adapter parameters and their associated candidate vectors, thereby improving the overall efficiency of the training process.
  • We experimentally validate SwitchLoRA on various sizes of the LLaMA model. SwitchLoRA shows significant perplexity improvements when compared to ReLoRA [34] and GaLore [60]. For the 1.3B model, SwitchLoRA achieves a perplexity of 15.01, surpassing the 15.23 perplexity obtained with full-rank training. Furthermore, by performing full fine-tuning on the resulting 1.3B model using the GLUE [48] tasks to validate the reasoning capabilities, we demonstrate that SwitchLoRA enhances model accuracy by approximately 1% on average, compared to the full-rank training method.
@@ 2 Methodology
A substantial body of research, such as various pruning methods [15, 4], has demonstrated that neural networks tend to exhibit low-rank characteristics after certain stages of training. Techniques for parameter-efficient fine-tuning, such as LoRA, capitalize on this observation. Concurrently, studies like [33, 13] have revealed that overparameterization in neural networks can lead to implicit regularization, thereby enhancing generalization. These findings underscore the importance of training with full parameters during the initial phase. Further empirical evidence supporting this phenomenon is provided in works like [49, 50, 34, 60]. Based on these insights, this section proposes a method designed to train a substantial number of parameters while selectively updating only a portion of the parameters at any one time to reduce communication overhead.
2.1 Low-Rank Adaptation (LoRA)
@@
Consider a pre-trained model with a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ from a specific linear layer. LoRA proposes an innovative modification: transforming $\mathbf{W}$ into $\mathbf{W}+\frac{\alpha}{r}\mathbf{B}\mathbf{A}$. Here, $\mathbf{B}\in\mathbb{R}^{m\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times n}$ are newly introduced matrices, where $r$ is a positive integer significantly smaller than both $m$ and $n$. And $\alpha$ is a constant hyperparameter, set to $r$ in the following description to clarify the algorithm’s mechanics. Then during fine-tuning, $\mathbf{W}$ is kept frozen while matrices $\mathbf{B}$ and $\mathbf{A}$ are trained. At the inference stage, $\mathbf{B}\mathbf{A}$ is added to $\mathbf{W}$ which preserves the model’s original structure. The matrix $\mathbf{A}$ is initialized using Kaiming initialization [17], while $\mathbf{B}$ is initially set to a zero matrix to ensure consistency.
2.2 SwitchLoRA
Below, we detail our proposed SwitchLoRA algorithm, the steps of which are outlined in Algorithm 1 and Algorithm 2.
Switching process
Now, let us delve deeper into the linear system $(\mathbf{W}+\mathbf{B}\mathbf{A})\mathbf{x}=\mathbf{y}$. As illustrated in Figure 1, we decompose the matrix $\mathbf{B}$ into its column vectors $\mathbf{b}_{k}\in\mathbb{R}^{m\times 1}$ for $k=1,\ldots,r$, represented as $\mathbf{B}=[\mathbf{b}_{1},\ldots,\mathbf{b}_{r}]$. Similarly, we decompose matrix $\mathbf{A}$ into its row vectors $\mathbf{a}_{k}^{T}\in\mathbb{R}^{1\times n}$ for $k=1,\ldots,r$, leading to $\mathbf{A}^{T}=[\mathbf{a}_{1}^{T},\ldots,\mathbf{a}_{r}^{T}]$. Hereafter, we call these vectors $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ as LoRA vectors.
@@ 2.2 SwitchLoRA
It is known for $k$ matrices $\mathbf{W}_{1},\mathbf{W}_{2},\ldots,\mathbf{W}_{k}$, the following inequality holds:
@@
If we adopt the strategy in LoRA to add $\mathbf{B}\mathbf{A}$ to $\mathbf{W}$ and only update $\mathbf{B}$ and $\mathbf{A}$ from the pre-training stage, according to (2), the rank of updated parameters of the local linear system through the entire training process will be limited to $2r$. This limitation can potentially impede the training efficacy. To mitigate this issue, we alter the values of $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ to $\mathbf{b}^{\prime}_{k}\in\mathcal{C}(\mathbf{B})$ and $\mathbf{a}^{\prime}_{k}\in\mathcal{C}(\mathbf{A}^{T})$ at appropriate frequencies, respectively, with these new values randomly selected from predefined candidate vectors list $\mathcal{C}(\mathbf{B})$ and $\mathcal{C}(\mathbf{A}^{T})$(one of $\mathbf{b}_{k}^{\prime}$ or $\mathbf{a}_{k}^{\prime}$ can be the same as $\mathbf{b}_{k}$ or $\mathbf{a}_{k}$). To maintain the consistency of the model’s output, we adjust $\mathbf{W}$ by adding the difference between the old and new LoRA components. To be more precise, when $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ are updated to $\mathbf{b}^{\prime}_{k}$ and $\mathbf{a}^{\prime}_{k}$, we accordingly adjust $\mathbf{W}$ with the equation $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{b}_{k}\mathbf{a}_{k}^{T}-\mathbf{b}_{k}^{\prime}\mathbf{a}_{k}^{\prime T}$.
@@
When implementing these updates, the updated parameters of both $\mathbf{B}$ and $\mathbf{A}$ are derived from $\min(m,n)$ distinct candidate vectors, which ensures updated parameters are full-rank. Readers can refer to [34, 62, 54] for more details.
When selecting candidate vectors, we have the option to choose randomly from $C(\mathbf{B})$ or $C(\mathbf{A}^{T})$. Alternatively, we can select candidate vectors sequentially from $C(\mathbf{B})$ or $C(\mathbf{A}^{T})$, restarting from the beginning once the end of the set is reached. We find that varying the matching orders of vectors $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ yields only minor differences in outcomes. A theoretical explanation for this phenomenon is provided in Appendix A. Additionally, to conserve GPU memory, spare candidate vectors can be offloaded to the CPU.
Switching frequency
@@
As mentioned in [11, 49, 34], the model initially exhibits full internal rank during pre-training, and the internal rank of each layer decreases progressively over time. Consequently, we have adopted an exponential decay function for the switching frequency, namely $frequency=Ce^{-\theta step}$, where the coefficients are determined empirically. Besides, the selection of LoRA rank $r$ for $\mathbf{B}\mathbf{A}$ is influenced by the final internal rank of the layers, which has been extensively explored in [21, 46, 57].
Optimizer states resetting
@@
Currently Large Language Models (LLMs) predominantly utilize Adam [26] and AdamW [36] optimizers over SGD, which rely on optimizer states. It is crucial to note that after switching LoRA vectors, the gradients associated with these parameters are also changed, which prevents the reuse of optimizer states. To address this issue, when $\mathbf{a}_{k}$ is switched, we reset the optimizer states of $\mathbf{b}_{k}$. And conversely, when $\mathbf{b}_{k}$ is switched, we reset optimizer states of $\mathbf{a}_{k}$. Note that we reset optimizer states of counterpart pair rather than optimizer states of the switched parameters itself. This approach will be further explained in Appendix A. Additionally, when the optimizer states are reset to zero, we freeze corresponding parameters for $N$ steps to maintain the robustness of the training. In this study, $N$ is set to $5$.
@@
Algorithm 2 SwitchLoRA training process. switch_num$(step,r,interval_{0},\theta)$ is an integer generator function which yields $\lfloor s\rfloor+X$ numbers sampled from $1$ to $r$ where $s=r/(interval_{0}e^{\theta step})$ and random variable $X\sim\text{Bernoulli}(s-\lfloor s\rfloor)$, i.e. $P(X=1)=1-P(X=0)=s-\lfloor s\rfloor$.
@@
4:        Freeze $\mathbf{W}$
5:        for $i$ in switch_num$(step,r,interval_{0},\theta)$ do
6:           Sample $j\sim\{k\}_{k=1}^{\min(m,n)}$
7:           $\mathbf{W},\mathbf{B},\mathbf{A}\leftarrow\text{switch}(\mathbf{W},\mathbf{B},\mathbf{A},i,j)$
8:           Freeze $\mathbf{A}_{i,:}$ for $N$ steps
9:        end for
10:        for $i$ in switch_num$(step,r,interval_{0},\theta)$ do
11:           Sample $j\sim\{k\}_{k=1}^{\min(m,n)}$
12:           $\mathbf{W}^{T},\mathbf{A}^{T},\mathbf{B}^{T}\leftarrow\text{switch}(\mathbf{W}^{T},\mathbf{A}^{T},\mathbf{B}^{T},i,j)$
13:           Freeze $\mathbf{B}_{:,i}$ for $N$ steps
@@
16:  end for
Initialization of SwitchLoRA
Results in [16, 56] have demonstrated the importance of initialization of LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ to the training effects. Unlike these works, which are applied only during the fine-tuning stage, our method is utilized throughout the entire training process. To achieve appropriate initialization for matrices $\mathbf{B}$ and $\mathbf{A}$ along with their candidate vectors, we follow the idea of Xavier initialization [12] and Kaiming initialization [17]. The derived initialization values are presented as follows: